OpenStudy (anonymous):

You have two block of clay in cube form and the edges are 10 cm. How many spheres with a radius of 5 cm can you make with that amount of clay?

OpenStudy (campbell_st):

Volume of a cube is V = 10^3 = 1000cm^3 2 cubes have a volume of 2000 cm^3 Volume of a sphere is \[V = \frac{4}{3} \pi r^3\] or in your question \[V = \frac{4}{3} \pi 5^3\] Once you have the volume of 1 cube with radius 5 cm divide that into the volume of the 2 cubes \[Number of Spheres = \frac{2000}{\frac{4}{3}\pi 5^3}\]

OpenStudy (campbell_st):

oops one you have the volume of 1 sphere divide into the volume of the 2 cubes

OpenStudy (anonymous):

Im not sure if my answer is right but i have "3 Sphere" but thank you @campbell_st for explaining it..

OpenStudy (campbell_st):

thats the answer I got... 3.8... which means only 3